Determine the regression line for the round-trip flight data. Predicted Cost-. +. Miles (Round to two decimal places as needed.) OA ??. 2000 4000 Miles 2000 4000 Miles 2000 4000 Miles Miles Interpret the slope and the intercept in the context of the problem. For every additional mile, on average, the price goes up by dollars. A trip of zero miles would cost about dollars. However, a trip would never be exactly zero miles, so this cost problem. (Round to two decimal places as needed.) sense in the context of the Using the regression line, how much would it cost, on average, to fly 500 miles? It would cost, on average, about to fly 500 miles. (Round to the nearest cent as needed.)The accompanying table shows the? round-trip fare for flights from one city to 10 other cities on major airlines. These were the lowest prices found online in early 2010. The airlines? varied, but the travel dates were all the same. How much would it? cost, on? average, to fly 500? miles? To answer this? question, perform a complete regression? analysis, including a scatterplot with a regression line.
Answer #1

(х-х) 183.000 2093519.6100 470458.8100 394.000 2094098.4100 84739.2100 317.000 1753240.8100 1502830.8100 374421.6100 448.000


predicted cost =164.88+0.08*miles

for every additional mile.........goes up by 0.08 dollars..................wuld cost about 164.88

dollars . However............so this cost does not make sense in ........

for x=500 ; predicted cost =164.88+0.08*500=204.88